# Learning an Optimal Assortment Policy under Observational Data

### v2

@@ 1 Introduction
Based on these results, we reveal another interesting finding: there exists a $\sqrt{K}$-statistical gap between uniform and non-uniform reward settings. In summary, we have provided a complete characterization of the statistical limit of offline assortment optimization under the MNL model, and presented a nearly minimax optimal algorithm, PRB, for this problem.
@@ 2 Related Work
However, most of these results focus on the dynamic learning setting, with Dong et al. (2023) as a exception. Specifically, the work by Dong et al. (2023) presents the only study on offline assortment optimization problems, where their algorithm relies on the optimal assortment coverage condition on data. In contrast, our algorithm requires only a much weaker optimal item coverage, and we prove this condition to be minimal.
@@ 4.1 Pessimistic Rank-Breaking
4:     $\widehat{p}_{j}\leftarrow\tau_{j}/\tau_{j0}$ //Compute the empirical probability that item $j$ beats 00
@@ 4.2 Theoretical Results
under our notation. This result involves the term $n_{S^{\star}}^{-1/2}$ with $n_{S^{\star}}:=\sum_{k=1}^{n}\bm{1}\{S_{k}=S^{\star}\}.$ It is straightforward to show that $\max_{i\in S^{\star}}n_{i}\geq n_{S^{\star}}$, and it is even possible for $\max_{i\in S^{\star}}n_{i}$ to be large while $n_{S^{\star}}=0$. Our result demonstrates that, instead of requiring observations of exact optimal assortments, it suffices to observe assortments that cover each item in the optimal assortment to learn the optimal policy. Beyond the improved from $n_{S^{\star}}$ to $\min_{i\in S^{\star}}n_{i}$, the problem parameter $\sqrt{N/\min_{i\in[N],\lvert S\rvert=K}p_{i}(S)}$ involved in (4.1) can blow up to +$\infty$+ ∞ as some $v_{i}$ approaches 00. Even in the best case, where all $p_{i}(S)=1/K$,this problem-dependent parameter scales as $\sqrt{NK}$, strictly larger than our result, which is solely linear in $K$.
@@ 5.1 Construction of Hard Problem Instances
    Significantly sub-optimal items $\mathcal{N}_{0}:=[N]\setminus[4K]$: each item in this subset has an attraction value of $1$ and a revenue of 00.
@@
By our construction, for every $(\bm{r},\bm{v})\in\mathcal{V}$, each item in $\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cup\mathcal{N}_{c}(\bm{r},\bm{v})$ is covered by $\{S_{1},\dots,S_{n}\}$ exactly $n_{\min}$ times. In each observation, only one of these items is accompanied by $K-1$ significantly sub-optimal items. In the following proof, we index each $S_{(\ell-1)n_{\min}+j}$ differently as $S_{j}^{(\ell)}$ for $\ell\in[4K]$ and $j\in[n_{\min}]$ to simplify our notation and emphasize that it is the $j$-th assortment containing item $\ell$.
@@ 5.3 Upper Bounding the KL-Divergence
$D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})$ $=D\left(\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}}(S_{j}^{(\ell)})\bigg{\lVert}\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}^{\prime}}(S_{j}^{(\ell)})\right)=n_{\min}\sum_{\ell=1}^{4K}D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1})).$
and the following result on each $D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1}))$:
@@ 6.2.1 Dependency on $\min_{i\in S^{\star}}n_{i}$
$r_{i}\equiv 1,\quad\forall i\in[N],\qquad\quad v_{i}=\begin{cases}\frac{1}{K}+\epsilon,&1\leq i\leq K,\\ \frac{1}{K},&i>K,\\ \end{cases}$
@@
the assortment generated by drawing $\{m_{1}^{(\ell)},\dots,m_{K-1}^{(\ell)}\}$ uniformly from all possible $K-1$ sized subsets of $[N]\setminus[4K].$ Given $S_{\ell}^{(j)}$, we let $i_{\ell}^{(j)}$ be the choice sampled from the MNL choice model with $\bm{v}$ condition on $S_{\ell}^{(j)}$. With this notation, we let the observed dataset $D_{\text{partial}}=\{i_{\ell}^{(j)},S_{\ell}^{(j)}\}_{j=1,\ell=1}^{n,2K}$. It can be seen from our construction that the optimal assortment is given by $S^{\star}=\{1,\dots,K\}$ and every item in $S^{\star}$ is covered by the observed assortments exactly for $n$ times.
@@ 6.2.2 Dependency on $K$ in uniform and non-uniform reward setting
$r_{i}\equiv 1,\quad\forall i\in[N],\qquad\quad v_{i}=\begin{cases}\frac{1}{K}+\epsilon,&1\leq i\leq K,\\ \frac{1}{K},&i>K,\\ \end{cases}$
@@ Appendix A Proof of Upper Bound Results
[…]
@@ Appendix B Proof of Auxiliary Results in Section 5
where we have used $\log(1+x)\leq x$ for $x>-1$ in the first inequality and $\sum_{j\in(S^{(\ell)}_{1})_{+}}(p_{j}-q_{j})=0$ in the second one.
Now if $\ell\notin(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cup\mathcal{N}_{\text{opt}}(\bm{r}^{\prime},\bm{v}^{\prime}))\setminus(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cap\mathcal{N}_{\text{opt}}(\bm{r}^{\prime},\bm{v}^{\prime}))$ we have then $p_{i}=q_{i},\quad\forall i\in(S^{(\ell)}_{1})_{+},$ that leads to the second inequality.
@@ C.1 Modifications on the hard-instance construction
    For items in $\mathcal{N}_{0}$, we modify the attraction value to $\frac{1}{K}$ and revenue to $1$
@@ C.2 Upper bounds of KL-divergence
$D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})$ $=D\left(\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}}(S_{j}^{(\ell)})\bigg{\lVert}\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}^{\prime}}(S_{j}^{(\ell)})\right)=n_{\min}\sum_{\ell=1}^{4K}D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1})).$
@@
$D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1}))\leq\begin{cases}5K(\epsilon^{\prime})^{2},&\text{if }\ell\in(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cup\mathcal{N}_{\text{opt}}(\bm{r}^{\prime},\bm{v}^{\prime}))\setminus(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cap\mathcal{N}_{\text{opt}}(\bm{r}^{\prime},\bm{v}^{\prime})),\\ 0,&\text{otherwise.}\end{cases}$ (C.2)
@@
Now if $\ell\notin(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cup\mathcal{N}_{\text{opt}}(\bm{r}^{\prime},\bm{v}^{\prime}))\setminus(\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cap\mathcal{N}_{\text{opt}}(\bm{r}^{\prime},\bm{v}^{\prime}))$ we have then $p_{i}=q_{i},\forall i\in(S^{(\ell)}_{1})_{+},$ that leads to the second inequality.

### v4

@@ 1 Introduction
Based on these results, we reveal another interesting finding: there exists a $\sqrt{K}$-statistical gap between uniform and non-uniform reward settings. In summary, we have provided a complete characterization of the statistical limit of offline assortment optimization under the MNL model, and presented a nearly minimax optimal algorithm, PRB, for this problem. In addition, our results provide three critical managerial insights:
  1. 1.
    We characterize in terms of a simple and interpretable metric the key characteristic (called the item coverage number) that defines whether existing historical data is necessary and sufficient for effective assortment optimization. Specifically, the item coverage number reflects the amount of information about each item in the optimal assortment that is contained within the offline dataset, capturing the essential data requirement for efficient offline assortment learning (see Table 1).
  2. 2.
    We provide a novel algorithm, called PRB, which does not require as input the item coverage number and yet achieves near-optimal performance (in terms of sub-optimality gap) by matching the theoretical minimax lower bounds for both uniform and non-uniform reward settings. PRB therefore guarantees that retailers can achieve the best possible results within the statistical limits of their available data.
  3. 3.
    Our research demonstrates that retailers only need to ensure each potentially optimal item appears individually within their historical data, eliminating the need to observe complete optimal assortment combinations. This finding dramatically simplifies data collection requirements, as it removes the impractical burden of gathering data on all possible assortment permutations, which grows exponentially with the number of products.
@@ 2 Related Work
However, most of these results focus on the dynamic learning setting, with Dong et al. (2023) as an exception. Specifically, the work by Dong et al. (2023) presents the only study on offline assortment optimization problems, where their algorithm relies on the optimal assortment coverage condition on data. In contrast, our algorithm requires only a much weaker optimal item coverage, and we prove this condition to be minimal.
@@ 4.1 Pessimistic Rank-Breaking
4:  $\widehat{p}_{j}\leftarrow\tau_{j}/\tau_{j0}$ //Compute the empirical probability that item $j$ beats 0
@@ 4.2 Theoretical Results
under our notation. This result involves the term $n_{S^{\star}}^{-1/2}$ with $n_{S^{\star}}:=\sum_{k=1}^{n}\bm{1}\{S_{k}=S^{\star}\}.$ It is straightforward to show that $\max_{i\in S^{\star}}n_{i}\geq n_{S^{\star}}$, and it is even possible for $\max_{i\in S^{\star}}n_{i}$ to be large while $n_{S^{\star}}=0$. Our result demonstrates that, instead of requiring observations of exact optimal assortments, it suffices to observe assortments that cover each item in the optimal assortment to learn the optimal policy. Beyond the improved from $n_{S^{\star}}$ to $\min_{i\in S^{\star}}n_{i}$, the problem parameter $\sqrt{N/\min_{i\in[N],\lvert S\rvert=K}p_{i}(S)}$ involved in (4.1) can blow up to $\infty$ as some $v_{i}$ approaches 0. Even in the best case, where all $p_{i}(S)=1/K$,this problem-dependent parameter scales as $\sqrt{NK}$, strictly larger than our result, which is solely linear in $K$.
@@ 5.1 Construction of Hard Problem Instances
    Significantly sub-optimal items $\mathcal{N}_{0}:=[N]\setminus[4K]$: each item in this subset has an attraction value of $1$ and a revenue of 0.
@@
By our construction, for every $(\bm{r},\bm{v})\in\mathcal{V}$, each item in $\mathcal{N}_{\text{opt}}(\bm{r},\bm{v})\cup\mathcal{N}_{c}(\bm{r},\bm{v})$ is covered by $\{S_{1},\dots,S_{n}\}$ exactly $n_{\min}$ times. In each observation, only one of these items is accompanied by K1K-1 significantly sub-optimal items. In the following proof, we index each $S_{(\ell-1)n_{\min}+j}$ differently as $S_{j}^{(\ell)}$ for $\ell\in[4K]$ and $j\in[n_{\min}]$ to simplify our notation and emphasize that it is the $j$-th assortment containing item $\ell$.
@@ 5.3 Upper Bounding the KL-Divergence
$D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})$ $).\displaystyle=D\left(\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}}(S_{j}^{(\ell)})\bigg{\lVert}\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}^{\prime}}(S_{j}^{(\ell)})\right)=n_{\min}\sum_{\ell=1}^{4K}D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1}$)).
and the following result on each $)D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1}$)):
@@ 6.2.1 Dependency on $\min_{i\in S^{\star}}n_{i}$
[…]
@@
the assortment generated by drawing $\{m_{1}^{(\ell)},\dots,m_{K-1}^{(\ell)}\}$ uniformly from all possible K1K-1 sized subsets of $N]\setminus[4K].$ Given $S_{\ell}^{(j)}$, we let $i_{\ell}^{(j)}$ be the choice sampled from the MNL choice model with $\bm{v}$ condition on $S_{\ell}^{(j)}$. With this notation, we let the observed dataset $D_{\text{partial}}=\{i_{\ell}^{(j)},S_{\ell}^{(j)}\}_{j=1,\ell=1}^{n,2K}$. It can be seen from our construction that the optimal assortment is given by $S^{\star}=\{1,\dots,K\}$ and every item in $S^{\star}$ is covered by the observed assortments exactly for $n$ times.
@@ 6.2.2 Dependency on $K$ in uniform and non-uniform reward setting
[…]
@@ 7 Conclusion
Acknowledgments.
This work is generously supported by the National Science Foundation under the grants CCF-2106508, CCF-2312204 and CCF-2312205.
@@ Appendix A Proof of Upper Bound Results
$\begin{aligned} &\sum_{k=1}^{n_{i}}\frac{1}{1+\sum_{j\in S_{k}}v_{j}}\geq 256\log(N/\delta)\\ &\qquad\implies 2\max\left\{2\sqrt{\sum_{k=1}^{n_{i}}\frac{(1+v_{i})}{1+\sum_{j\in S_{k}}v_{j}}},\sqrt{\log(1/\delta^{\prime})}\right\}=4\sqrt{\sum_{k=1}^{n_{i}}\frac{(1+v_{i})}{1+\sum_{j\in S_{k}}v_{j}}}\\ &\qquad\implies 4\sqrt{\sum_{k=1}^{n_{i}}\frac{(1+v_{i})}{1+\sum_{j\in S_{k}}v_{j}}\cdot 2\log(N/\delta)}\leq\frac{1}{2}\sum_{k=1}^{n_{i}}\frac{1+v_{i}}{1+\sum_{j\in S_{k}}v_{j}}\\ &\qquad\implies\tau_{i0}\geq\frac{1}{2}\sum_{k=1}^{n_{i}}\frac{1+v_{i}}{1+\sum_{j\in S_{k}}v_{j}}\geq 69\log(N/\delta)\end{aligned}$ (A.2)
@@ Appendix B Proof of Auxiliary Results in Section 5
where we have used $\log(1+x)\leq x$ for x>1x>-1 in the first inequality and $\sum_{j\in(S^{(\ell)}_{1})_{+}}(p_{j}-q_{j})=0$ in the second one.
[…]
@@ C.1 Modifications on the hard-instance construction
    For items in $\mathcal{N}_{0}$, we modify the attraction value to $\frac{1}{K}$ and revenue to $1.$
@@ C.2 Upper bounds of KL-divergence
$D(\mathbb{P}_{\bm{r},\bm{v}}\lVert\mathbb{P}_{\bm{r}^{\prime},\bm{v}^{\prime}})$ $).\displaystyle=D\left(\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}}(S_{j}^{(\ell)})\bigg{\lVert}\prod_{\ell=1}^{4K}\prod_{j=1}^{n_{\min}}\mathbb{Q}_{\bm{v}^{\prime}}(S_{j}^{(\ell)})\right)=n_{\min}\sum_{\ell=1}^{4K}D(\mathbb{Q}_{\bm{v}}(S^{(\ell)}_{1})\rVert\mathbb{Q}_{\bm{v}^{\prime}}(S^{(\ell)}_{1}$)).
@@
[…]
@@
[…]